A local FM radio station broadcasts at a frequency of 106.0 MHz.

Calculate the energy of the frequency at which it is broadcasting.
Energy = ___ kJ/photon

(1 MHz = 106 sec -1)
Answer #1

Energy of the given frequency can be calculated by the equation

E = h

E = Energy

h = Planck's Constant

h =  6.626 * 10-34 J.s

= Frequency

= 106 MHZ   (1MHZ = 1 * 106 HZ)

E = 6.626 * 10-34 J.s ( 106 * 106 s-1 )

As s and s-1 gets cancelled,

E = 7.024 * 10-21 KJ/Photon

Energy of Frequency at which the local FM radio station broadcasts is 7.024 * 10-21 KJ/Photon